OpenStudy (anonymous):

8x^1/2<---- is it 4x yes or no ?

OpenStudy (ryan123345):

no

OpenStudy (anonymous):

oh

OpenStudy (ryan123345):

anything to the 1/2 power is basically the same thing as square rooting it.

OpenStudy (anonymous):

hmmm....

OpenStudy (ryan123345):

c: yup

OpenStudy (anonymous):

but square root 8 is 2 rad 2

OpenStudy (anonymous):

is it \(8x^{\frac{1}{2}}\) or \((8x)^{\frac{1}{2}}\) ?

OpenStudy (ryan123345):



OpenStudy (anonymous):

well the equation is 8x^1/2=16

OpenStudy (anonymous):

@satelite 2nd one

OpenStudy (anonymous):

\[\sqrt{8x^2}=\sqrt{8}\times\sqrt{x^2}=\sqrt{2^2\times2}\times x=2x\sqrt{2}\]

OpenStudy (ryan123345):

yea that is a good point @satellite73 because if it is 8x^1/2 then you would find the square root of x and multiply 8. If it was (8x)^1/2 then you would find the square root of the whole thing (8x)